# MIGHTY: Hermite Spline-based Efficient Trajectory Planning

### v2

@@ IV-B Representation Benchmarking: Complex Case
TABLE II: Complex-scene benchmark across 24 goals and five speed limits ($v_{\max}\!\in\!\{1,2,3,4,5\}$ m/s). GC = GCOPTER, MI = MIGHTY.
@@
We further evaluate MIGHTY in a complex environment (Fig. 3). The start is fixed at $(0,0,0.5)\,$\mathrm{m}$\$. For each $v_{\max}\in\{1,2,3,4,5\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, we plan to 24 goal locations with $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=2.5\,$\mathrm{m}$\$. Optimization settings are identical to the simple case for both methods. We report the following metrics: $v_{\max}$ [m/s]: Speed limit, $T_{\mathrm{comp}}$ [ms]: Computation time, $T_{\mathrm{trav}}$ [s]: Travel time, $L_{\mathrm{path}}$ [m]: Path length, $\int\lVert\mathbf{j}(t)\rVert\,dt$ [m/s2]: Jerk smoothness integral, and $\rho_{\mathrm{viol}}$ [%]: Fraction of time violating dynamic bounds.
@@
IV-C Ablation Study on Scaled and Unscaled Variables
@@
As discussed in Section III-D, we optimize over scaled variables to improve numerical stability. We evaluate this with an ablation between MIGHTY (scaled) and an unscaled variant. As in the complex benchmarking, both run in the same environment with 24 goals for $v_{\max}\in\{1,\dots,5\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, using identical objectives, constraints, initial guess, and solver settings. Fig. 4 shows that the scaled variant is about $2\times$ faster in computation time, with lower jerk, while path length and travel time are nearly identical. These results confirm that scaling improves numerical stability and optimization efficiency without degrading trajectory quality.
IV-D Benchmarking in Static Environments
@@
TABLE III: Benchmarking results (safety, computation time, performance, and constraint violation). We mark in green the best value in each column considering only configurations with $R_{\mathrm{succ}}=100\%$; configurations with $R_{\mathrm{succ}}<100\%$ are excluded from best highlighting. For EGO-Swarm2, we evaluated both its default sensor, depth camera, and a LiDAR (as used by SUPER and MIGHTY) and report results as “camera”  |  “LiDAR”. We also tested EGO-Swarm with LiDAR, but its performance was substantially worse, so we omit it from the table for clarity. For SUPER, computation time is reported as “exploratory”  &  “safe”. The data are shaded red if they have unsuccessful rates $R_{\mathrm{succ}}<100\%$.
@@
To thoroughly evaluate MIGHTY, we benchmark it in an obstacle-rich environment (Fig. 5) against EGO-Swarm [zhou2021ego-swarm], EGO-Swarm2 [zhou2022swarm], and SUPER [ren2025super]. Static cylindrical obstacles are placed randomly with radii sampled in $1.0\text{\,}\mathrm{m}$,$1.5\text{\,}\mathrm{m}$\$\,] and height $6.0\text{\,}\mathrm{m}$, occupying a $300\text{\,}\mathrm{m}$\times$40\text{\,}\mathrm{m}$$ area. The agent starts at $(0,0,3)\,$\mathrm{m}$\$ and the goal is $(305,0,3)\,$\mathrm{m}$\$. We constrain velocity, acceleration, and jerk to $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$10.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$. For global guidance we run JPS [harabor2011online], then build a safe-flight corridor (SFC) [liu2017planning] for local optimization.
@@
$J=$ $~w_{T}\!\sum_{s}T_{s}+w_{\text{smooth}}\tilde{J}_{\text{smooth}}+w_{\text{SFC}}J_{\text{SFC}}$
@@
where $J_{\text{SFC}}\;=\;\sum_{s}\int_{0}^{T_{s}}\sum_{h=1}^{H_{s}}\phi_{\mu}\!\big(\,\mathbf{a}_{s,h}^{\top}\mathbf{x}(t)-b_{s,h}+C_{\text{SFC}}\,\big)\,dt,$ penalizes leaving the SFC, $J_{v}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{v}(t)\|^{2}-v_{\max}^{2})\,dt$, $J_{a}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{a}(t)\|^{2}-a_{\max}^{2})\,dt$, and $J_{j}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{j}(t)\|^{2}-j_{\max}^{2})\,dt$, Note that the jerk smoothness cost $\tilde{J}_{\text{smooth}}$ uses the closed form given in Eq. 6, which avoids per-sample recomputation. MIGHTY’s weights are $w_{T}=$5\text{\times}{10}^{2}$\$, $w_{\text{smooth}}=${10}^{-1}$\$, $w_{\text{SFC}}=${10}^{3}$\$, $w_{v}=${10}^{3}$\$, $w_{a}=${10}^{3}$\$, $w_{j}=${10}^{3}$\$. We run 10 trials per method. Metrics: $R_{\mathrm{succ}}$ [%] (success rate; collision-free, reaches goal); $T_{\mathrm{opt}}$ [ms] (local optimization time); $T_{\mathrm{total}}$ [ms] (total planning time); $T_{\mathrm{trav}}$ [s] (travel time); $L_{\mathrm{path}}$ [m] (path length); $S_{\mathrm{jerk}}=\!\int\!\lVert\mathbf{j}(t)\rVert\,dt$ [m/s3] (smoothness); $\bar{S}_{\mathrm{jerk}}=\sqrt{\tfrac{1}{T}\int\!\lVert\mathbf{j}(t)\rVert^{2}\,dt}$ [m/s3] (RMS jerk; time-normalized); $\rho_{\mathrm{vel}},\,\rho_{\mathrm{acc}},\,\rho_{\mathrm{jerk}}$ [%] (velocity/acceleration/jerk violations). We report $\bar{S}_{\mathrm{jerk}}$ to account for differing travel times across methods.
@@ IV-E Dynamic Environments
TABLE IV: Nearest dynamic-obstacle distances by environment for 10 runs each. Columns show 0th (min), 5th, 50th (median), and 95th percentiles. The robot collision radius is set to $0.1\text{\,}\mathrm{m}$.
Environment Distance [m]
0th (min) 5th 50th (median) 95th
Only Dynamic 0.8 2.1 3.2 5.2
Dynamic + Static 1.0 2.5 3.8 6.8
We further evaluate MIGHTY in two dynamic scenarios: (1) only dynamic obstacles, and (2) dynamic & static obstacles. For both, in addition to the cost terms described in Section IV-D, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$. We set the weight $w_{\text{dyn}}=${10}^{1}$\$ and use a soft barrier radius $C_{\text{dyn}}=$3.0\text{\,}\mathrm{m}$\$. Note that $C_{\text{dyn}}$ is where the penalty activates, not a hard safety distance. The robot collision radius is $0.1\text{\,}\mathrm{m}$. The start and goal are $(0,0,3)\,$\mathrm{m}$\$ and $(100,0,3)\,$\mathrm{m}$\$, and, in this more complex environment, the dynamic constraints are set to $v_{\max}=$2.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$.
Dynamic-only environment: We generate 100 dynamic obstacles following smooth trefoil-knot trajectories, with initial positions $x\in[0,100]\,$\mathrm{m}$\$, $y\in[-10,10]\,$\mathrm{m}$\$, $z\in[1,6]\,$\mathrm{m}$\$. Fig. 6 shows an example of the environment and the trajectory. MIGHTY avoids all obstacles in 10/10 trials. As shown in Table IV, the minimum nearest-obstacle distance is $0.8\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$ safety threshold.
Dynamic & static environment: We add 50 dynamic obstacles (same motion model) to the static forest as shown in Fig. 7 and run 10 trials. MIGHTY avoids all obstacles in all trials. Table IV summarizes nearest-obstacle distances, where the minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$.
@@ V Hardware Experiments
We also evaluated MIGHTY in three hardware experiment: (1) long-duration flights at $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, (2) high-speed flights up to $v_{\max}=7\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, and (3) flights with obstacles introduced during the mission. The perception uses a Livox Mid-360 LiDAR, and localization is provided by DLIO [chen2023dlio]. Planning runs onboard on an Intel NUC 13, and low-level control uses PX4 [meier2015px4] on a Pixhawk flight controller [meier2011pixhawk]. All perception, planning, and control run in real time onboard.
Figure 8: Long-duration experiment with $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$. Top: partial action sequence at $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$ (cropped for clarity). Bottom: full trajectory histories overlaid on the recorded LiDAR point cloud. Warmer colors indicate higher speed (red: $v_{\max}$, blue: 0.0 $0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$); white points are the onboard point cloud.
@@
Figure 9: High-speed experiments at $v_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$6.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. Each panel shows the action sequence with trajectory history over the recorded point cloud. Warmer colors indicate higher speed (red: $v_{\max}$, blue: 0.0 $0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$); white points are the onboard point cloud.
@@
Dynamic Obstacle Flight Experiment: To assess robustness to changes in the environment, we introduce and move obstacles during the mission (Fig. 10). We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Section IV-E for details. The agent successfully avoids all obstacles and reaches the goal without collision over $490\text{\,}\mathrm{s}$ of flight time.
@@ VI Conclusions
VII Acknowledgements
The authors would like to thank Lili Sun for her help with the hardware experiments and Mason B. Peterson for his support with robotdatapy library. The authors also thank Juan Rached for his help with hardware setup. This work was supported in part by DSTA and ARL grant W911NF-21-2-0150.

### v3

@@ IV-D Representation Benchmarking: Complex Case
TABLE III: Complex-scene benchmark across 24 goals and five speed limits ($v_{\max}\!\in\!\{1,2,3,4,5\}$ m/s). GC = GCOPTER, MI = MIGHTY.
@@
We further evaluate MIGHTY in a complex environment (Fig. 5) with start at $(0,0,0.5)\,$\mathrm{m}$\$ and 24 goals at $x,y\in[-15,15]\,$\mathrm{m}$\$ and $z=2.5\,$\mathrm{m}$\$. Optimization settings are identical to the simple case for both methods. We report the following metrics: $v_{\max}$ [m/s]: Speed limit, $T_{\mathrm{comp}}$ [ms]: Computation time, $T_{\mathrm{trav}}$ [s]: Travel time, $L_{\mathrm{path}}$ [m]: Path length, $\int\lVert\mathbf{j}(t)\rVert\,dt$ [m/s2]: Jerk smoothness integral, and $\rho_{\mathrm{viol}}$ [%]: Fraction of time violating dynamic bounds.
@@
IV-E Ablation Study on Scaled and Unscaled Variables
@@
As discussed in Section III-D, we optimize over scaled variables to improve numerical stability. We evaluate this via ablation using the same complex benchmarking setup. Fig. 6 shows the scaled variant is $\approx 2\times$ faster with lower jerk, while path length and travel time are nearly identical, confirming that scaling improves efficiency without degrading quality.
IV-F Benchmarking in Static Environments
@@
TABLE IV: Benchmarking results (safety, computation time, performance, and constraint violation). We mark in green the best value in each column considering only configurations with $R_{\mathrm{succ}}=100\%$; configurations with $R_{\mathrm{succ}}<100\%$ are excluded from best highlighting. For EGO-Swarm2, we evaluated both its default sensor, depth camera, and a LiDAR (as used by SUPER and MIGHTY) and report results as “camera”  |  “LiDAR”. We also tested EGO-Swarm with LiDAR, but its performance was substantially worse, so we omit it from the table for clarity. For SUPER, computation time is reported as “exploratory”  &  “safe”. The data are shaded red if they have unsuccessful rates $R_{\mathrm{succ}}<100\%$.
@@
To thoroughly evaluate MIGHTY, we benchmark it in an obstacle-rich environment (Fig. 7) against EGO-Swarm [27], EGO-Swarm2 [26], and SUPER [16]. Static cylindrical obstacles are placed randomly with radii sampled in $1.0\text{\,}\mathrm{m}$,$1.5\text{\,}\mathrm{m}$\$\,] and height $6.0\text{\,}\mathrm{m}$, occupying a $300\text{\,}\mathrm{m}$\times$40\text{\,}\mathrm{m}$$ area. The agent starts at $(0,0,3)\,$\mathrm{m}$\$ and the goal is $(305,0,3)\,$\mathrm{m}$\$. We constrain velocity, acceleration, and jerk to $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$10.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$. For global planning we run A*, then build a safe-flight corridor (SFC) [10] for local optimization. The intermediate waypoints for A* are projected from the terminal goal onto the local occupancy map centered on the drone. The average computation time for A* and SFC generation is $0.26\text{\,}\mathrm{m}\mathrm{s}$ and $0.48\text{\,}\mathrm{m}\mathrm{s}$, respectively.
@@
$J=$ $~w_{T}\!\sum_{s}T_{s}+w_{\mathrm{smooth}}\tilde{J}_{\text{smooth}}+w_{\text{SFC}}J_{\text{SFC}}$
@@
where $J_{\text{SFC}}\;=\;\sum_{s}\int_{0}^{T_{s}}\sum_{h=1}^{H_{s}}\phi_{\mu}\!\big(\,\mathbf{a}_{s,h}^{\top}\mathbf{x}(t)-b_{s,h}+C_{\text{SFC}}\,\big)\,dt,$ penalizes leaving the SFC, $J_{v}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{v}(t)\|^{2}-v_{\max}^{2})\,dt$, $J_{a}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{a}(t)\|^{2}-a_{\max}^{2})\,dt$, and $J_{j}=\sum_{s}\!\int_{0}^{T_{s}}\phi_{\mu}(\|\mathbf{j}(t)\|^{2}-j_{\max}^{2})\,dt$, Note that the jerk smoothness cost $\tilde{J}_{\text{smooth}}$ uses the closed form given in Eq. 6, which avoids per-sample recomputation. MIGHTY’s weights are $w_{T}=$5\text{\times}{10}^{2}$\$, $w_{\mathrm{smooth}}=${10}^{-1}$\$, $w_{\text{SFC}}=${10}^{3}$\$, $w_{v}=${10}^{3}$\$, $w_{a}=${10}^{3}$\$, $w_{j}=${10}^{3}$\$. We run 10 trials per method. Metrics: $R_{\mathrm{succ}}$ [%] (success rate; collision-free, reaches goal); $T_{\mathrm{opt}}$ [ms] (local optimization time); $T_{\mathrm{total}}$ [ms] (total planning time); $T_{\mathrm{trav}}$ [s] (travel time); $L_{\mathrm{path}}$ [m] (path length); $S_{\mathrm{jerk}}=\!\int\!\lVert\mathbf{j}(t)\rVert\,dt$ [m/s3] (smoothness); $\bar{S}_{\mathrm{jerk}}=\sqrt{\tfrac{1}{T}\int\!\lVert\mathbf{j}(t)\rVert^{2}\,dt}$ [m/s3] (RMS jerk; time-normalized); $\rho_{\mathrm{vel}},\,\rho_{\mathrm{acc}},\,\rho_{\mathrm{jerk}}$ [%] (velocity/acceleration/jerk violations). We report $\bar{S}_{\mathrm{jerk}}$ to account for differing travel times across methods.
@@ IV-G Dynamic Environments
We further evaluate MIGHTY in two dynamic scenarios: (1) only dynamic obstacles, and (2) dynamic & static obstacles. For both, in addition to the cost terms described in Section IV-F, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$. We set the weight $w_{\text{dyn}}=${10}^{1}$\$ and use a soft barrier radius $C_{\text{dyn}}=$3.0\text{\,}\mathrm{m}$\$. Note that $C_{\text{dyn}}$ is where the penalty activates, not a hard safety distance. The robot collision radius is $0.1\text{\,}\mathrm{m}$. The start and goal are $(0,0,3)\,$\mathrm{m}$\$ and $(100,0,3)\,$\mathrm{m}$\$, and, in this more complex environment, the dynamic constraints are set to $v_{\max}=$2.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$.
Dynamic-only environment: We generate 100 dynamic obstacles following smooth trefoil-knot trajectories, with initial positions $x\in[0,100]\,$\mathrm{m}$\$, $y\in[-10,10]\,$\mathrm{m}$\$, $z\in[1,6]\,$\mathrm{m}$\$. Fig. 8 shows an example of the environment and the trajectory. MIGHTY avoids all obstacles in 10/10 trials. The minimum nearest-obstacle distance is $0.8\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$ safety threshold.
Dynamic & static environment: We add 50 dynamic obstacles (same motion model) to the static forest as shown in Fig. 9 and run 10 trials. MIGHTY avoids all obstacles in all trials. The minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$.
@@ V Hardware Experiments
We also evaluated MIGHTY in three hardware experiment: (1) long-duration flights at $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, (2) high-speed flights up to $v_{\max}=7\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, and (3) flights with obstacles introduced during the mission. The perception uses a Livox Mid-360 LiDAR, and localization is provided by DLIO [2]. Planning runs onboard on an Intel NUC 13, and low-level control uses PX4 [11] on a Pixhawk flight controller [12]. All perception, planning, and control run in real time onboard.
Figure 10: Long-duration experiment with $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$. Top: action sequence at $v_{\max}=$4.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. Bottom: trajectory histories overlaid on the recorded LiDAR point cloud. Warmer colors indicate higher speed (red: $v_{\max}$, blue: 0.0 $0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$); white points are the onboard point cloud.
@@
Figure 11: High-speed experiments at $v_{\max}=$5.0\text{\,}$,$6.0\text{\,}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. Each panel shows the action sequence with trajectory history over the recorded point cloud.
@@
Dynamic Obstacle Flight Experiment: To assess robustness to changes in the environment, we introduce and move obstacles during the mission (Fig. 12). We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Section IV-G for details. The agent successfully reaches the goal without collision over $490\text{\,}\mathrm{s}$.